OpenStudy (anonymous):

f(x)=ax+b given that f(2)=1 and f(5)=7 solve equation of ff(X)

OpenStudy (bahrom7893):

f(2)=1, and f(5)=7, so we can build two equations with unknowns a and b

OpenStudy (bahrom7893):

1 = a2+b 7 = a5+b

OpenStudy (bahrom7893):

or if i write it properly: 2a+b=1 5a+b=7

OpenStudy (anonymous):

alright i have a =2.5 and b=-4

OpenStudy (bahrom7893):

b = 1-2a 5a+1-2a=7 3a = 6 a=2

OpenStudy (bahrom7893):

b = 1-4=-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now lets solve

OpenStudy (bahrom7893):

that is, we've solved it. f(x) = 2x-3

OpenStudy (anonymous):

no its ff(x)

OpenStudy (anonymous):

we should solve for f^2

OpenStudy (anonymous):

which means we are going to substitute our values using the values of a and b on the equation

OpenStudy (bahrom7893):

ohhh then just multiply it by itself.. i think

OpenStudy (anonymous):

Um getting a different answer i just don know y the correct answer is 2.25

OpenStudy (bahrom7893):

i honestly dont even understand the question

OpenStudy (anonymous):

Oh soorry it says the equation ffx=0

OpenStudy (bahrom7893):

what's ff(x)?

OpenStudy (bahrom7893):

ohh wait

OpenStudy (bahrom7893):

f(f(X)) = 2(2x-3)-3 = 0 4x - 6 - 3 = 0 4x = 9 x = 9/4

OpenStudy (anonymous):

I was reading from one textbook it means f^2

OpenStudy (bahrom7893):

No it doesn't mean f^2.. well not really

OpenStudy (anonymous):

wow explain

OpenStudy (bahrom7893):

f(x) = 2x-3, right? or f(a) = 2a-3 or f(random) = 2 random - 3?

OpenStudy (anonymous):

yeah ryt

OpenStudy (bahrom7893):

do u understand what I'm doing there? \[f(\Delta)=2\Delta-3\]

OpenStudy (bahrom7893):

So now you have: f(f(x)), but f(x) = 2x-3, so replace the x in the f(x) with 2x-3

OpenStudy (bahrom7893):